3 moles of A and 4 moles of B are mixed together and allowed to come into equilibrium according to the following reaction.
            3A(g) + 4B(g) = 2C(g) +3D(g)
 when equilibrium is reached, There is 1 mole of C. The equilibrium constant of the reaction is 
  1.   1÷4
  2.    1÷3
  3.     1÷2
  4.      1
​Please explain in deail...............
 

Grade:12th pass

2 Answers

Arun
5 years ago
t
Dear student
 
I think that you have typed a wrong question. Please check all the reaction as I am not getting answer  as the options are.
 
You can also attachan image. I will be happy to help you.
 
Arun
5 years ago
Dear student
 
it has been solved, all thing is correct
 
solution
 
 
3A + 4 B = 2C + 3 D
 
at equilibrium
3- 3x  4- 4x    2x     3x
 
as 2x = 1
x = ½
 
now equilibrium constant = 6x^2 / (3- 3x) (4-4x) = 6/4 / 3 = 2/4 = ½
 
hence option C is correct